The angular momentum of a flywheel having a rotational inertia of 0.125 kgm2decreasesfrom 3.0 to 2.0 kgm2/s in a period of 1.5 s. (a) What is the average torque acting on the flywheelduring this period? (b) Assuming a uniform angular acceleration, through how many revolutions will the flywheel have turned? (c) How much work was done on the wheel? (d) What was theaverage power supplied by the flywheel?
1
Expert's answer
2020-11-17T07:10:04-0500

"T_{avg}=\frac{L_f-L_i}{\delta t}"


a) "\frac {3.0-2.0}{1.5}= 1.15N.m"


b) "\theta= \frac{(3)(1.5s)+(-1.467) \frac{(1.5)^{2}}{2}}{0.125}"


"\theta=22.8 rad"


c)"W=\tau\theta= (-1.47)(22.8)"


"W=-33.51J"


"P_{avg}=\frac {-33.51}{1.50}= 22.3W"